OpenStudy (anonymous):

A social club charters a bus at a cost of $2805 to take a group of members on an excursion to Atlantic City. At the last minute, 4 people in the group decide not to go. This raises the transportation cost per person by $4. How many people originally intended to take the trip?

OpenStudy (anonymous):

so how many people were originally intended to take the trip?

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

its a long solution and so had to send in two parts....☺

OpenStudy (anonymous):

i appreciate it ugh.. i chose the wrong version its supposed to be A social club charters a bus at a cost of $7350 to take a group of members on an excursion to Atlantic City. At the last minute, 1 people in the group decide not to go. This raises the transportation cost per person by $3. How many people originally intended to take the trip

OpenStudy (anonymous):

the numbers all switched. is there anyway for you to figure that out quick

OpenStudy (anonymous):

well... let me see.....

OpenStudy (anonymous):

so once again lol how many were orginally intended

OpenStudy (anonymous):

wait more coming...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

?

OpenStudy (anonymous):

geez crazy equation

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry friend.... i seem to have made some mistakes in the solution... give me your email id and ill send the solution in a few minutes for both questions.....

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

i just need to know how many people orginally intended to take the trip

OpenStudy (anonymous):

just send it on here or too gfoste00@leeu.edu if thats easier

OpenStudy (anonymous):

then b patient, i'll post the solution in five minutes...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

are you posting it on he?

OpenStudy (anonymous):

here*

OpenStudy (anonymous):

yes, I am first posting for the first question which I had solved wrong.....

OpenStudy (anonymous):

let the total number of people originally = x let the money paid by each person = y x*y = 2805 -------------- (1) 2805 y = -------- -----------(2) x when 4 member do not go, then members going = x-4 now money these (x-4) members have to pay = (y + 4) (x-4)(y+4) = 2805 xy - 4y + 4x - 16 = 2805 Substituting value of xy from (1) 2805 – 4y + 4x – 16 = 2805 2789 – 4y + 4x = 2805 -4y + 4x = 2805 – 2789 -4y + 4x = 16 4(-y + x) = 16 -y + x = 16/4 Substituting value of y from (2)

OpenStudy (anonymous):

-2805 ------ + x = 4 x -2805 + x² ---------- = 4 x -2805 + x² = 4x x² - 4x - 2805 = 0 x² - 55x + 51x - 2805 = 0 x(x-55) +51(x-55)=0 (x - 55)(x + 51) = 0 if x - 55 = 0 => x = 55 if x+51 = 0 => x = -51 (we ignore this value as number of people cannot be negative) so number of people originally taking the trip was 55

OpenStudy (anonymous):

now I'll post for the second one in a couple of minutes.....

OpenStudy (anonymous):

so the final answer is?

OpenStudy (anonymous):

wait this isnt even the right one. i need the solution for the one with 7805

OpenStudy (anonymous):

this one : A social club charters a bus at a cost of $7350 to take a group of members on an excursion to Atlantic City. At the last minute, 1 people in the group decide not to go. This raises the transportation cost per person by $3. How many people originally intended to take the trip

OpenStudy (anonymous):

let the total number of people originally = x let the money paid by each person = y x*y = 7350 -------------- (1) 7350 y = -------- -----------(2) x when 1 member does not go, then members going = x-1 now money these (x-1) members have to pay = (y + 3) (x-1)(y+3) = 7350 xy - y + 3x - 3 = 7350 Substituting value of xy from (1) 7350 – 3 -y + 3x = 7350 - y + 3x = 7350 - 7347 - y + 3x = 3

OpenStudy (anonymous):

more follows...

OpenStudy (anonymous):

Substituting value of y from (2) - 7350 ------ + 3x = 3 x -7350 + 3x² ----------- = 3 x -7350 + 3x² = 3x 3x² - 3x - 7350 = 0 3[x² - x - 2450] = 0 3[x² -50x + 49x - 2450] = 0 3[x(x - 50) +49(x - 50)] = 0 3(x - 50)(x + 49) = 0

OpenStudy (anonymous):

if x-50 = 0, then x = 50 if x+49 = 0,then x = -49 (we ignore this value as number of people cannot be negative) so the number of people originally going on the trip was 50